OpenStudy (anonymous):

[6.05] Your piggy bank has a total of 20 coins in it; some are dimes and some are nickels. If you have a total of $1.25, how many nickels do you have?

OpenStudy (texaschic101):

d = dimes n = nickels d + n = 20 --> d = 20 - n .10d + .05n = 125 sub 20 - n in for d in the 2nd equation and solve for n . Once you find n sub it back into the 1st equation to find d. You can also check your answers by subbing them both into the original equations and if they come out equal, then you have done the problem correctly.